Mike's age is three fourths of Mary's age. In 5 years Mike's age will be seven ninths Mary's age. How old is Mike now? Please help with this problem.
Answer by checkley77(12844)

LET X=MARY'S AGE NOW.
LET 3X/4=MIKE'S AGE NOW.
IN 5 YEARS MARY WILL BE X+5
IN 5 YEARS MIKE WILL BE (3X/4+5)
(3X/4+5)=7/9(X+5)
(3X+20)/4=(7X+35)/9
9(3X+20)=4(7X+35)
27X+180=28X+140
27X-28X=140-180
-X=-40
X=40 WHICH IS MARY'S AGE NOW
3/4*40=30 IS MIKE'S AGE NOW
PROOF
30+5=7/9(40+5)
35=7/9*45
35=315/9
35=35